OpenStudy (anonymous):

what is the slope of 9x-3y=23?

OpenStudy (anonymous):

y = mx + b The slope is m. b is the y-intercept.

OpenStudy (nirmalnema):

-A/B a=coeff...of x b = coeff...of y so slope is equal to - 9/(-3) =3

OpenStudy (anonymous):

thanks!